# Eco-evolutionary dynamics of cooperative antimicrobial resistance in a population of fluctuating volume and size

### v2

@@ Abstract
Antimicrobial resistance to drugs (AMR), a global threat to human and animal health, is often regarded as resulting from a cooperative behaviour. Moreover, microbes generally evolve in volatile environments that, together with demographic fluctuations (birth and death events), drastically alter population size and strain survival. Motivated by the need to better understand the evolution of AMR, we study a population of time-varying size consisting of two competing strains, one drug-resistant and one drug-sensitive, subject to demographic and environmental variability. This is modelled by a binary carrying capacity randomly switching between mild and harsh environmental conditions, and driving the fluctuating volume (total amount of nutrients and antimicrobials at fixed concentration), and thus the size of the community (number of resistant and sensitive cells). We assume that AMR is a shared public good when the concentration of resistant cells exceeds a fixed concentration cooperation threshold, above which the sensitive strain has a growth advantage, whereas resistant cells dominate below it. Using computational means, and devising an analytical treatment (built on suitable quenched and annealed averaging procedures), we fully characterise the influence of fluctuations on the eco-evolutionary dynamics of AMR, and notably obtain specific strain fixation and long-lasting coexistence probabilities as a function of the environmental variation rate and cooperation threshold. We find that microbial strains tend to coexistence, but demographic fluctuations eventually lead to the extinction of resistant or sensitive cells for small or large values of the concentration cooperation threshold, respectively. This also holds for dynamic environments, whose specific properties determine the extinction timescale.
@@ 1 Introduction
AMR, whose rise is a major societal threat [50, 51], can often be interpreted as resulting from a cooperative behaviour. This occurs when antimicrobial drugs are inactivated by an enzyme produced, at a metabolic cost, by resistant cells [52, 53, 54, 55, 56, 13]. When the concentration of resistant microbes exceeds a fixed “concentration cooperation threshold” there are enough drug-inactivating enzymes for the protection against the toxin to be shared with sensitive cells at no metabolic cost. However, when the concentration of resistant cells is below the cooperation threshold, only resistant microbes benefit from the protection of the enzymes. In this scenario, AMR is mediated by drug-inactivating enzymes that act as public good only above the cooperation threshold. This yields the spread of the resistant strain below the threshold, where the drug limits the growth of (or kills) sensitive microbes, while these thrive above the cooperative threshold where the drug-inactivating enzyme concentration is high [52, 53, 54, 13]. In a large population and static environment, sensitive and resistant strains thus coexist [54, 55, 57, 58]. However, this picture can be greatly altered by fluctuations arising in finite population subject to EV [30, 27, 28, 29, 59, 60, 42, 13].
@@ 2.1 Model
We consider a well-mixed population of time-varying size $N(t)=N_{R}(t)+N_{S}(t)$ consisting of $N_{R}$ resistant ($R$) microbes and $N_{S}$ sensitive ($S$) microbes, which compete for the same resources in the presence of a constant influx of antimicrobial drug. The strain $R$ experiences a fixed metabolic cost (reduced reproduction rate) to constantly produce resistance enzymes [^1] As opposed to the case where $R$ can regulate the production of resistance enzyme by quorum sensing [64, 65]. The influence of this mechanism on the evolution of AMR is not addressed in this work. that break down the drug in their immediate vicinity, protecting them and also reducing the overall drug concentration in the community. Microbes of strain $S$ pay no metabolic cost associated to drug-inactivating enzymes (of which they are no producers), but their growth is hampered by the presence of the toxin. When the fraction of $R$ in the population is sufficiently high, they produce enough resistance enzymes to bring the overall drug concentration below its minimum inhibitory concentration (MIC) [52, 66, 54, 58]. When this happens, the drug is ineffective everywhere and $S$ benefits from the enzymes protection despite not having to pay any cost for their production. The production of the resistance enzymes by $R$ making up a sufficiently high fraction of the community can hence be regarded as cooperative behaviour, and their protection against antimicrobials as a form of public good [54, 55, 56], see figure 1(a).
For concreteness, we assume a biostatic antimicrobial that only affects the growth rate of $S$ [^2] The case where the drug increases the death rate of the $S$ microbes corresponds to a biocidal antimicrobial, and is not directly considered here. This is a stringent limitation, since a given drug may act as a biocidal or biostatic antimicrobial depending on its concentration [67]. The metabolic cost for $R$ to produce the enzymes is denoted by $s$ and the growth hindrance experienced by $S$ microbes in the presence of the drug (above the MIC) is denoted by $a$. In this setting, the fraction of $R$ in the population, given by $x\equiv N_{R}/N$, determines whether the enzymes offer a public good protection to $S$. If $x_{\text{th}}$ is the fraction of $R$ necessary for the enzymes to share their protection across the community, $S$ is protected when $x\geq x_{\text{th}}$, whereas it is affected by the drug when $x<x_{\text{th}}$. Accordingly, the growth fitness $f_{R/S}$ of the $R$ and $S$ microbes are
@@ 2.2 Environmental fluctuations & master equation
This reflects the fact that here variations of $K(t)$ are accompanied by those of the volume (amount of nutrients and toxins) at fixed concentration, see figure 1(b).
@@
Figure 1: Model. (a) Top: when the fraction of $R$ (blue microbes) is below the concentration cooperation threshold $x_{\text{th}}$, antimicrobial drug hinders the growth rate of $S$ (red microbes) and $R$ has then a growth advantage. Bottom: in a cooperative scenario arising above the concentration cooperation threshold, resistance becomes shared (green shields) as the fraction of $R$ exceeds $x_{\text{th}}$ and these generate enough resistance enzymes (public good, green shade) to break down the drug and set its concentration below the MIC for the whole community. (b) Ecological dynamics as random extensive changes. The environment has a constant concentration of nutrients and antimicrobial drug, but switches from high to low volume (i.e. from high to low carrying capacity) and vice-versa at rates $\nu_{+}$ and $\nu_{-}$, respectively. The number of microbes of each species, $N_{R}$ and $N_{S}$, evolves in this changing environment according to equation (2.2). (c) Temporal eco-evolutionay dynamics of the microbial community for parameters $x_{\text{th}}=0.37$, $s=0.1$, $a=0.25$, $K_{-}=100$, $K_{+}=1000$, $\nu=0.1$, $\delta=0.5$, and for initial conditions $K(t=0)=K_{+}$, $N_{R}(t=0)=x_{\text{th}}K_{+}$, and $N_{S}(t=0)=\left(1-x_{\text{th}}\right)K_{+}$; thick black line shows the sample path of the time-switching carrying capacity $K(t)$, thick solid coloured lines correspond to a realisation of the piecewise deterministic Markov process that ignores demographic fluctuations and is defined by equations (8) and (14) (see section 4.2), for the total population ($N$, green), number of $R$ ($N_{R}=xN$, blue), and number of $S$ ($N_{S}=(1-x)N$, red); noisy lines are the corresponding stochastic realisation of the full model under the joint effect of demographic and environmental fluctuations; dashed lines show the piecewise (meta-)stable equilibrium $N_{R}=x_{\text{th}}K(t)$ (blue) and $N_{R}=(1-x_{\text{th}})K(t)$ (red). (d) $R$ fraction $x=N_{R}/N$ for the same sample path of varying environment as in (c); line styles as in panel (c). See text for more details.

### v3

@@ Abstract
Antimicrobial resistance to drugs (AMR), a global threat to human and animal health, is often regarded as resulting from cooperative behaviour. Moreover, microbes generally evolve in volatile environments that, together with demographic fluctuations (birth and death events), drastically alter population size and strain survival. Motivated by the need to better understand the evolution of AMR, we study a population of time-varying size consisting of two competing strains, one drug-resistant and one drug-sensitive, subject to demographic and environmental variability. This is modelled by a binary carrying capacity randomly switching between mild and harsh environmental conditions, and driving the fluctuating volume (total amount of nutrients and antimicrobials at fixed concentration), and thus the size of the community (number of resistant and sensitive cells). We assume that AMR is a shared public good when the concentration of resistant cells exceeds a fixed concentration cooperation threshold, above which the sensitive strain has a growth advantage, whereas resistant cells dominate below it. Using computational means, and devising an analytical treatment (built on suitable quenched and annealed averaging procedures), we fully characterise the influence of fluctuations on the eco-evolutionary dynamics of AMR, and notably obtain specific strain fixation and long-lasting coexistence probabilities as a function of the environmental variation rate and cooperation threshold. We find that microbial strains tend to coexistence, but demographic fluctuations eventually lead to the extinction of resistant or sensitive cells for small or large values of the concentration cooperation threshold, respectively. This also holds for dynamic environments, whose specific properties determine the extinction timescale.
@@ 1 Introduction
AMR, whose rise is a major societal threat [50, 51], can often be interpreted as resulting from a cooperative behaviour. This occurs when antimicrobial drugs are inactivated by an enzyme produced, at a metabolic cost, by resistant cells [52, 53, 54, 55, 56, 13]. When the concentration of resistant microbes exceeds a fixed “concentration cooperation threshold” there are enough drug-inactivating enzymes for the protection against the toxin to be shared with sensitive cells at no metabolic cost. However, when the concentration of resistant cells is below the cooperation threshold, only resistant microbes benefit from the protection of the enzymes. In this scenario, AMR is mediated by drug-inactivating enzymes that act as a public good only above the cooperation threshold. This yields the spread of the resistant strain below the threshold, where the drug limits the growth of (or kills) sensitive microbes, while these thrive above the cooperative threshold where the drug-inactivating enzyme concentration is high [52, 53, 54, 13]. In a large population and static environment, sensitive and resistant strains thus coexist [54, 55, 57, 58]. However, this picture can be greatly altered by fluctuations arising in finite population subject to EV [30, 27, 28, 29, 59, 60, 42, 13].
@@ 2.1 Model
We consider a well-mixed population of time-varying size $N(t)=N_{R}(t)+N_{S}(t)$ consisting of $N_{R}$ resistant ($R$) microbes and $N_{S}$ sensitive ($S$) microbes, which compete for the same resources in the presence of a constant influx of antimicrobial drug. The strain $R$ experiences a fixed metabolic cost (reduced reproduction rate) to constantly produce resistance enzymes [^1] As opposed to the case where $R$ can regulate the production of resistance enzyme by quorum sensing [64, 65]. The influence of this mechanism on the evolution of AMR is not addressed in this work. that break down the drug in their immediate vicinity, protecting them and also reducing the overall drug concentration in the community. Microbes of strain $S$ pay no metabolic cost associated to drug-inactivating enzymes (of which they are not producers), but their growth is hampered by the presence of the toxin. When the fraction of $R$ in the population is sufficiently high, they produce enough resistance enzymes to bring the overall drug concentration below its minimum inhibitory concentration (MIC) [52, 66, 54, 58]. When this happens, the drug is ineffective everywhere and $S$ benefits from the enzymes protection despite not having to pay any cost for their production. The production of the resistance enzymes by $R$ making up a sufficiently high fraction of the community can hence be regarded as cooperative behaviour, and their protection against antimicrobials as a form of public good [54, 55, 56], see figure 1(a).
For concreteness, we assume a biostatic antimicrobial that only affects the growth rate of $S$ [^2] The case where the drug increases the death rate of the $S$ microbes corresponds to a biocidal antimicrobial, and is not directly considered here. This is a stringent limitation, since a given drug may act as a biocidal or biostatic antimicrobial depending on its concentration [67]. The metabolic cost for $R$ to produce the enzymes is denoted by $s$ and the growth hindrance experienced by $S$ microbes in the presence of the drug (above the MIC) is denoted by $a$. In this setting, the fraction of $R$ in the population, given by $x\equiv N_{R}/N$, determines whether the enzymes offer a public good protection to $S$. If $x_{\text{th}}$ is the fraction of $R$ necessary for the enzymes to share their protection across the community, $S$ is protected when $x\geq x_{\text{th}}$, whereas it is affected by the drug when $x<x_{\text{th}}$. Accordingly, the growth fitnesses $f_{R/S}$ of the $R$ and $S$ microbes are
@@ 2.2 Environmental fluctuations & master equation
This reflects the fact that variations of $K(t)$ are accompanied by those of the volume (amount of nutrients and toxins) at fixed concentration, see figure 1(b).
@@
Figure 1: Model. (a) Top: when the fraction of $R$ (blue microbes) is below the concentration cooperation threshold $x_{\text{th}}$, antimicrobial drug hinders the growth rate of $S$ (red microbes) and $R$ has a growth advantage. Bottom: in a cooperative scenario arising above the concentration cooperation threshold, resistance becomes shared (green shields) as the fraction of $R$ exceeds $x_{\text{th}}$ and these generate enough resistance enzymes (public good, green shade) to break down the drug and set its concentration below the MIC for the whole community. (b) Ecological dynamics as random extensive changes. The environment has a constant concentration of nutrients and antimicrobial drug, but switches from high to low volume (i.e. from high to low carrying capacity) and vice-versa at rates $\nu_{+}$ and $\nu_{-}$, respectively. The number of microbes of each species, $N_{R}$ and $N_{S}$, evolves in this changing environment according to equation (2.2). (c) Temporal eco-evolutionary dynamics of the microbial community for parameters $x_{\text{th}}=0.37$, $s=0.1$, $a=0.25$, $K_{-}=100$, $K_{+}=1000$, $\nu=0.1$, $\delta=0.5$, and for initial conditions $K(t=0)=K_{+}$, $N_{R}(t=0)=x_{\text{th}}K_{+}$, and $N_{S}(t=0)=\left(1-x_{\text{th}}\right)K_{+}$; thick black line shows the sample path of the time-switching carrying capacity $K(t)$, thick solid coloured lines correspond to a realisation of the piecewise deterministic Markov process that ignores demographic fluctuations and is defined by equations (8) and (14) (see section 4.2), for the total population ($N$, green), number of $R$ ($N_{R}=xN$, blue), and number of $S$ ($N_{S}=(1-x)N$, red); noisy lines are the corresponding stochastic realisation of the full model under the joint effect of demographic and environmental fluctuations; dashed lines show the piecewise (meta-)stable equilibrium $N_{R}=x_{\text{th}}K(t)$ (blue) and $N_{R}=(1-x_{\text{th}})K(t)$ (red). (d) $R$ fraction $x=N_{R}/N$ for the same sample path of varying environment as in (c); line styles as in panel (c). See text for more details.